# Beamfocusing Optimization for Near-Field Wideband Multi-User Communications

### v4

@@ V Numerical Results
In this section, numerical results obtained through Monte Carlo simulations are provided to verify the effectiveness of the proposed FAD and HTS approaches. The following simulation setup is exploited throughout our simulations unless otherwise specified. It is assumed that the BS is equipped with $N=512$ antennas with half-wavelength spacing, $N_{\mathrm{RF}}=4$ RF chains, and $N_{\mathrm{T}}=16$ TTDs for each RF chain. The maximum time delay of each TTD is set to $t_{\max}=N/(2f_{c})=1.28$ nanosecond (ns) [15]. The central frequency, system bandwidth, number of subcarriers, and length of the cyclic prefix are assumed to be $f_{c}=100$ GHz, $B=10$ GHz, $M=10$, and $L_{\mathrm{CP}}=4$, respectively. There are $K=4$ near-field communication users randomly located in close proximity to the BS from $5$ m to $15$ m. The number of NLoS channels between the BS and each user is assumed to be $L_{k}=4,\forall k$. The average reflection coefficient of NLoS channels is set to $-15$ dB. The noise power density is set to $-174$ dBm/Hz. For the proposed algorithms, the convergence threshold is set to $10^{-3}$. The penalty factors are set to $\rho=\bar{\rho}=10^{3}$, and their reduction factor is set to $c=0.5$. Since the performance of Algorithms 2 and 3 is highly sensitive to the initialized parameters, the optimization variables of these two algorithms are initialized based on the HTS approach to guarantee a good performance. All following results are obtained by averaging over $100$ random channel realizations. For performance comparison, we consider the following two benchmark schemes:
  • Optimal Digital beamforming (BF): In this benchmark, each antenna is connected to a dedicated RF chain, which realizes the full-dimensional frequency-dependent beamforming. This benchmark provides an upper-bound performance. The fully-digital beamformers can be optimized using the methods in [29, 31, 32].
  • Conventional BF: In this benchmark, the analog beamforming is achieved by only frequency-independent PSs. The corresponding beamforming optimization problem can be solved through the proposed penalty-FDA algorithm by setting $t_{n,l}=0$.
Figure 6: Average array gain over all frequencies and directions realized by different analog beamforming methods for a user $5$ m away from the BS.
V-A Evaluation of Proposed Heuristic Analog Beamforming
@@
For near-field analog beamforming design, there have been some other heuristic designs in the literature, such as the near-field phase-delayed focusing (PDF) method [20] and the far-field delayed phase precoding (DPP) method [13, 15]. The near-field PDF scheme is based on a piecewise-far-field approximation, which approximates the array response of each sub-array as a far-field channel, while the far-field DPP scheme approximates the array response of the whole array through a far-field channel. However, the proposed method relies on the exact near-field channel rather than the far-field approximation. Furthermore, in these existing methods, the maximum time delay constraint is not taken into account, which may lead to further performance loss. In Fig. 6, we compare the average array gain achieved by different analog beamforming methods. In this case, the conventional method uses only the PS to generate the beam, which is aligned to the user position at the center frequency. It can be seen that the proposed method achieves significant performance gains compared to the existing method due to the use of a more accurate approximation and the optimization of the maximum delay constraint.
@@
V-B Spectral Efficiency Versus Transmit Power
In Fig. 7, we evaluate the spectral efficiency achieved by the different schemes with respect to the transmit power $P_{t}$. It can be seen that the fully-connected TTD-BF architecture performs close to the optimal digital BF architecture by employing a limited number of TTDs. For the subconnected TTD-BF architecture, although only one sub-array per RF chain can be used to transmit the signal, the resulting performance loss is not significant. In addition, the FDA approach is the most efficient approach in both fully-connected and sub-connected architectures due to its ability to achieve a stationary point solution. Although only partial CSI is used to heuristically design the analog beamformer, the performance loss of the HTS method is minimal compared to the FDA method. On the contrary, conventional BF architectures utilizing only frequency-dependent PSs suffer from significant performance degradation due to significant spatial wideband effects.
V-C Spectral and Energy Efficiency Versus Number of TTDs
In Fig. 8, we investigate the effect of the number of TTDs on the spectral efficiency of the considered communication system. The results show that as the number of TTDs increases, the performance of the fully-connected architecture gradually converges towards the optimal digital BF performance. This improvement is attributed to the enhanced mitigation of spatial wideband effects by installing more TTDs. However, this improvement is not evident in the sub-connected architecture. In other words, the sub-connect architecture requires a much smaller number of TTDs to reach its optimal performance due to the weaker spatial wideband effect of the small-aperture sub-arrays connected to each RF chain. This result is consistent with the analysis in Remark 1. Furthermore, an intriguing phenomenon is that the HTS method in the fully-connected architecture experiences a significant performance degradation when the number of TTDs is limited (i.e., $N_{T}\leq 8$). This is because the piecewise-near-field approximation becomes invalid in this case, see Proposition 1. On the contrary, the performance of the HTS method is consistently comparable to that of the FDA method in the sub-connected architecture. This is also expected since the piecewise-near-field approximation is only needed to approximate the array response of small-aperture sub-arrays, which requires fewer TTDs to ensure high accuracy.
@@ V-C Spectral and Energy Efficiency Versus Number of TTDs
In Fig. 9, we further demonstrate the energy efficiency versus the number of TTDs. Energy efficiency is defined as the ratio between spectral efficiency and power consumption. In particular, let $P_{\mathrm{BB}}=300$ mW, $P_{\mathrm{RF}}=200$ mW, $P_{\mathrm{PS}}=30$ mW, and $P_{\mathrm{TTD}}=100$ mW denote the power consumption of baseband processing, a RF chain, a PS, and a TTD, respectively [15]. Then, the power consumption of optimal digital BF architectures is given by $P_{\mathrm{FD}}=P_{t}+P_{\mathrm{BB}}+NP_{\mathrm{RF}}$. The power consumption of the fully-connected and sub-connected TTD-BF architectures is given by $P_{f}=P_{t}+P_{\mathrm{BB}}+N_{\mathrm{RF}}P_{\mathrm{RF}}+NN_{\mathrm{RF}}P_{\mathrm{PS}}+NN_{\mathrm{T}}P_{\mathrm{TTD}}$ and $P_{s}=P_{t}+P_{\mathrm{BB}}+N_{\mathrm{RF}}P_{\mathrm{RF}}+NP_{\mathrm{PS}}+NN_{\mathrm{T}}P_{\mathrm{TTD}}$, respectively. Similarly, the power consumption of conventional BF architectures can be calculated by excluding the power consumption attributed to TTDs. As can be observed from Fig. 9, the sub-connected architecture is much more energy-efficient than the fully-connected architecture. Furthermore, it is evident that the energy efficiency of the fully-connected architecture has peak points at $N_{\mathrm{T}}=8$ and $16$. This is because, beyond this point, the additional increase in the number of TTDs leads to marginal performance improvements in spectral efficiency, as depicted in Fig. 8, while incurring a substantial power consumption. Conversely, the energy efficiency of the sub-connected architecture exhibits a continuous decline from the outset, as the installation of additional TTDs fails to provide significant improvements in spectral efficiency.
@@
V-D Spectral Efficiency Versus Maximum Delay of TTDs
V-E Spectral Efficiency Versus Bandwidth
In Fig. 11, the effect of the system bandwidth $B$ on the spectral efficiency performance is investigated. For a fair comparison, the transmit power $P_{t}$ is adjusted according to the bandwidth so that the transmit signal-to-noise ratio $P_{t}/\sigma_{m,k}^{2}$ remains constant. It can be seen that the performance of the fully-connected architecture decreases significantly as the bandwidth increases. This is because the larger the bandwidth, the more severe the spatial wideband effect is, and more TTDs are required to mitigate it effectively. In contrast, the sub-connected architecture exhibits only a slight performance loss with increasing bandwidth. This is due to the smaller aperture of the sub-arrays connected to each RF chain, which results in a slight spatial wideband effect even with larger bandwidths. Furthermore, since the accuracy of the piecewise near-field approximation depends on the number of TTDs and the bandwidth, see Proposition 1, the performance of the HTS method in fully-connected architectures exhibits a significant performance degradation as the bandwidth increases. However, in sub-connected architectures, the small size of the subarrays reduces the effect of large bandwidths, thus making the HTS method more efficient.
@@ VI Conclusion
This paper investigated the beamforming optimization problem in near-field wideband multi-user communication systems with the aim of mitigating the performance loss due to spatial wideband effects. A new sub-connected architecture was proposed for TTD-based hybrid beamforming architectures. A pair of wideband beamforming methods, i.e., FDA and low-complexity HTS methods, were proposed for both fully-connected and sub-connected architectures. Numerical results showed that although the sub-connected architecture leads to performance degradation in terms of spectral efficiency compared to the fully-connected architecture, it is much more energy-efficient and robust for the increase in the system bandwidth and the reduction in the maximum delay of the TTDs.
@@ Appendix A Proof of Proposition 1
$\vartheta_{l}=0,\quad\tilde{\nu}_{l,q}=\nu_{l}-\tilde{\chi}_{q}d,\quad f_{M}-f_{c}\approx\frac{B}{2},\quad\forall l,q.$ (70)
@@
$G(f_{M},0,r)\approx$ $\frac{1}{N}\left|\sum_{l=1}^{N_{\mathrm{T}}}\sum_{q=1}^{\tilde{N}}e^{-j\pi\eta B\tilde{\chi}_{q}}\right|=\left|\frac{N_{\mathrm{T}}}{N}\sum_{q=-\frac{\tilde{N}-1}{2}}^{\frac{\tilde{N}-1}{2}}e^{-j\pi\eta Bq}\right|$
$=$ $\left|\frac{N_{\mathrm{T}}}{N}\frac{\sin\left(\frac{\pi BN}{4f_{c}N_{\mathrm{T}}}\right)}{\sin\left(\frac{\pi B}{4f_{c}}\right)}\right|,$ (71)

### v6

@@ V Numerical Results
In this section, numerical results obtained through Monte Carlo simulations are provided to verify the effectiveness of the proposed FAD and HTS approaches. The following simulation setup is exploited throughout our simulations unless otherwise specified. It is assumed that the BS is equipped with $N=512$ antennas with half-wavelength spacing, $N_{\mathrm{RF}}=4$ RF chains, and $N_{\mathrm{T}}=16$ TTDs for each RF chain. The maximum time delay of each TTD is set to $t_{\max}=N/(2f_{c})=2.56$ nanosecond (ns) [16]. The central frequency, system bandwidth, number of subcarriers, and length of the cyclic prefix are assumed to be $f_{c}=100$ GHz, $B=10$ GHz, $M=10$, and $L_{\mathrm{CP}}=4$, respectively. There are $K=4$ near-field communication users randomly located in close proximity to the BS from $5$ m to $15$ m. The number of NLoS channels between the BS and each user is assumed to be $L_{k}=4,\forall k$. The average reflection coefficient of NLoS channels is set to $-15$ dB. The noise power density is set to $-174$ dBm/Hz.
For the proposed algorithms, the convergence threshold is set to $10^{-3}$. The penalty factors are set to $\rho=\bar{\rho}=10^{3}$, and their reduction factor is set to $c=0.5$. For the one-dimensional search, we set $Q=10^{3}$. Furthermore, “HTS-PNF” represents the HTS approach based on the PNF-based analog beamforming design, while “HTS-Robust” refers to the HTS approach using the robust design. Since the performance of the FDA approach might be sensitive to the initialized parameters, the optimization variables of these two algorithms are initialized based on the HTS-PNF approach to guarantee a good performance. All following results are obtained by averaging over $100$ random channel realizations. For performance comparison, we mainly consider the following two benchmark schemes: 1) Optimal Digital beamforming (BF): In this benchmark, each antenna is connected to a dedicated RF chain, which realizes the full-dimensional frequency-dependent beamforming. This benchmark provides an upper-bound performance. The fully-digital beamformers can be optimized using the methods in [34, 40, 41]; 2) Conventional BF: In this benchmark, the analog beamforming is achieved by only frequency-independent PSs. The corresponding beamforming optimization problem can be solved through the proposed penalty-FDA algorithm by setting $t_{n,l}=0,\forall n,l$.
Figure 6: Convergence behavior of the proposed FDA approaches (left) and robust design for analog beamforming optimization (right).
V-A Convergence and Complexity of Proposed Approaches
Fig. 6 demonstrates the convergence behavior of the proposed iterative algorithms. To quantitatively analyze the complexity, we examine the CPU time of each approach using MATLAB R2024a on an Apple M3 silicon chip. For the fully-connected architecture, the CPU times are 37.91 s for FDA, 0.53 s for HTS-Robust, and 0.35 s for HTS-PNF. For the sub-connected architecture, the CPU times are 18.52 s, 0.42 s, and 0.37 s, respectively. The HTS approach significantly reduces computational complexity by avoiding joint optimization of analog and digital beamformers. Additionally, the FDA approach for sub-connected architecture consumes less CPU time than the fully-connected architecture due to the simplified optimization process enabled by the spatial structure of the sub-connected analog beamformers.
V-B Evaluation of Proposed Heuristic Analog Beamforming
In Fig. 7, we evaluate the normalized array gain achieved by the proposed heuristic analog beamforming method across different frequencies and bandwidths. Here, we consider the fully-connected architecture and a single user located at $(45^{\circ},10\text{ m})$ as an example. We compare our approach against the following wideband analog beamforming designs for conventional BF architecture: 1) Mean channel covariance matrix (MCCM) method [13], which designs the analog beamformer based on the averaged covariance matrix of array response vectors over all subcarriers; 2) Mean channel matrix (MCM) method [13], which designs the analog beamformer based on the averaged array response vector over all subcarriers; and 3) Central frequency (CF) method, where the design is based on the array response at the central frequency. It can be observed that the proposed heuristic method employing TTDs significantly enhances the array gain throughout the entire frequency band. However, with bandwidth expansion from $10$ GHz to $30$ GHz, the proposed PNF-based method maintains high array gain only over a limited range of the frequency band. In contrast, the robust design consistently achieves high array gain over the entire band, confirming its robustness.
Figure 7: Normalized array gain across the frequency band with $10$ GHz (top), $20$ GHz (middle), and $30$ GHz (bottom) bandwidths.
@@ V-B Evaluation of Proposed Heuristic Analog Beamforming
In Fig. 9, we further evaluate the proposed heuristic analog beamforming method with different maximum time delays. In addition to the above benchmarks, we also consider two existing wideband analog beamforming methods using TTD for performance comparison, namely the near-field phase-delayed focusing (PDF) method [21] and the far-field delayed phase precoding (DPP) method [14, 16]. The near-field PDF scheme is based on a piecewise-far-field approximation, which approximates the array response of each sub-array as a far-field channel, while the far-field DPP scheme approximates the array response of the whole array through a far-field channel. It can be seen from Fig. 9 that the proposed method achieves significant performance gains compared to the existing method due to the use of a more accurate approximation and the optimization of the maximum delay constraint.
V-C Spectral Efficiency Versus Transmit Power
In Fig. 9, we evaluate the spectral efficiency achieved by the different schemes with respect to the transmit power $P_{t}$. It can be seen that the fully-connected TTD-BF architecture performs close to the optimal digital BF architecture by employing a limited number of TTDs. For the subconnected TTD-BF architecture, although only one sub-array per RF chain can be used to transmit the signal, the resulting performance loss is not significant. In addition, the FDA approach is the most efficient approach in both fully-connected and sub-connected architectures due to its ability to achieve a stationary point solution. Although only partial CSI is used to heuristically design the analog beamformer, the performance loss of the HTS approaches is negligible compared to the FDA approach. On the contrary, conventional BF architectures utilizing only frequency-dependent PSs suffer from significant performance degradation due to significant spatial wideband effects.
Figure 8: Average array gain versus maximum time delay, $t_{\max}$.
@@ V-C Spectral Efficiency Versus Transmit Power
V-D Spectral and Energy Efficiency Versus Number of TTDs
In Fig. 11, we investigate the effect of the number of TTDs on the spectral efficiency of the considered communication system. The results show that as the number of TTDs increases, the performance of the fully-connected architecture gradually converges towards the optimal digital BF performance. This improvement is attributed to the enhanced mitigation of spatial wideband effects by installing more TTDs. However, this improvement is not evident in the sub-connected architecture. In other words, the sub-connect architecture requires a much smaller number of TTDs to reach its optimal performance due to the weaker spatial wideband effect of the small-aperture sub-arrays connected to each RF chain. This result is consistent with the analysis in Remark 1. Furthermore, an intriguing phenomenon is that the HTS-PNF approach in the fully-connected architecture experiences a significant performance degradation when the number of TTDs is limited (i.e., $N_{T}\leq 8$). This is because the PNF approximation becomes invalid in this case, see Proposition 1. In this case, the HTS-Robust approach should be adopted to guarantee the performance. On the contrary, the performance of the HTS-PNF approach is consistently comparable to that of the FDA approach in the sub-connected architecture. This is also expected since the PNF approximation is only needed to approximate the array response of small sub-arrays, which requires fewer TTDs to ensure high accuracy.
In Fig. 11, we further demonstrate the energy efficiency versus the number of TTDs. Energy efficiency is defined as the ratio between spectral efficiency and power consumption. In particular, let $P_{\mathrm{BB}}=300$ mW, $P_{\mathrm{RF}}=200$ mW, $P_{\mathrm{PS}}=30$ mW, and $P_{\mathrm{TTD}}=100$ mW denote the power consumption of baseband processing, a RF chain, a PS, and a TTD, respectively [16]. Then, the power consumption of optimal digital BF architectures is given by $P_{\mathrm{FD}}=P_{t}+P_{\mathrm{BB}}+NP_{\mathrm{RF}}$. The power consumption of the fully-connected and sub-connected TTD-BF architectures is given by $P_{\mathrm{F\textendash TTD}}=P_{t}+P_{\mathrm{BB}}+N_{\mathrm{RF}}P_{\mathrm{RF}}+NN_{\mathrm{RF}}P_{\mathrm{PS}}+NN_{\mathrm{T}}P_{\mathrm{TTD}}$ and $P_{\mathrm{S\textendash TTD}}=P_{t}+P_{\mathrm{BB}}+N_{\mathrm{RF}}P_{\mathrm{RF}}+NP_{\mathrm{PS}}+NN_{\mathrm{T}}P_{\mathrm{TTD}}$, respectively. Similarly, the power consumption of conventional BF architectures can be calculated by excluding the power consumption attributed to TTDs. As can be observed from Fig. 11, the sub-connected architecture is much more energy-efficient than the fully-connected architecture. Furthermore, it is evident that the energy efficiency of the fully-connected architecture has peak points at $N_{\mathrm{T}}=8$ and $16$. This is because, beyond this point, the additional increase in the number of TTDs leads to marginal performance improvements in spectral efficiency, as depicted in Fig. 11, while incurring a substantial power consumption. Conversely, the energy efficiency of the sub-connected architecture exhibits a continuous decline from the outset, as the installation of additional TTDs fails to provide significant improvements in spectral efficiency.
@@ V-D Spectral and Energy Efficiency Versus Number of TTDs
V-E Spectral Efficiency Versus Maximum Delay of TTDs
@@
V-F Spectral Efficiency Versus Bandwidth
In Fig. 13, we study how the system bandwidth $B$ impacts spectral efficiency. We keep the transmit power $P_{t}$ adjusted with the bandwidth so that the transmit signal-to-noise ratio $P_{t}/\sigma_{m,k}^{2}$ remains consistent. It can be seen that as bandwidth increases, the performance of the fully-connected architecture declines significantly due to the more significant spatial wideband effect, requiring more TTDs to counter it effectively. In contrast, the sub-connected architecture shows negligible performance loss with increasing bandwidth because its smaller sub-array apertures result in a less severe spatial wideband effect. Furthermore, the HTS-PNF approach in fully-connected architectures experiences notable performance loss as bandwidth increases, which is primarily due to the reduced accuracy of the piecewise near-field approximation, as detailed in Proposition 1. Compared to the HTS-PNF approach, the HTS-Robust performs more robustly against the increased bandwidth in the fully-connected architectures. Conversely, in sub-connected architectures, the smaller size of the sub-arrays diminishes the impact of increased bandwidths, thereby enhancing the efficiency of the HTS-PNF approach.
@@ VI Conclusion
This study explored beamforming optimization in near-field wideband multi-user communication systems to counteract spatial wideband effects. We introduced a new sub-connected architecture for TTD-based hybrid beamforming and proposed two wideband beamforming methods, FDA and low-complexity HTS, applicable to both fully-connected and sub-connected architectures. Numerical results indicated that while the sub-connected architecture sacrifices some spectral efficiency, it significantly enhances energy efficiency and robustness against increases in system bandwidth and reductions in TTD delays.
@@ Appendix A Proof of Proposition 1
$\vartheta_{l}=0,\quad\tilde{\nu}_{l,q}=\nu_{l}-\tilde{\chi}_{q}d,\quad f_{M}-f_{c}\approx\frac{B}{2},\quad\forall l,q.\vspace{-0.2cm}$ (76)
@@
$G(f_{M},0,r)\approx\frac{1}{N}\left|\sum_{l=1}^{N_{\mathrm{T}}}\sum_{q=1}^{\tilde{N}}e^{-j\pi\eta B\tilde{\chi}_{q}}\right|$
$=\left|\frac{N_{\mathrm{T}}}{N}\sum_{q=-\frac{\tilde{N}-1}{2}}^{\frac{\tilde{N}-1}{2}}e^{-j\pi\eta Bq}\right|=\left|\frac{N_{\mathrm{T}}}{N}\frac{\sin\left(\frac{\pi BN}{4f_{c}N_{\mathrm{T}}}\right)}{\sin\left(\frac{\pi B}{4f_{c}}\right)}\right|,$ (77)
@@
Appendix B Proof of Proposition 2
For the objective function of problem (61), we get the following result from the triangle inequality:
$\left|\sum_{l=1}^{N_{\mathrm{T}}}e^{-j2\pi f_{m}\left(\frac{\nu_{l}(\theta_{k},r_{k})-r_{k}}{c}+t_{k,l}\right)}\right|\leq N_{\mathrm{T}},$ (78)
where the maximum value is achieved when all summation terms have the same arbitrary phase. We set this arbitrary phase to $-2\pi f_{m}\zeta_{k}$. The optimal $t_{k,l}$ must satisfy
$-2\pi f_{m}\left(\frac{\nu_{l}(\theta_{k},r_{k})-r_{k}}{c}+t_{k,l}\right)=-2\pi f_{m}\zeta_{k}$ (79)
[…]
To ensure that $t_{k,l}\geq 0,\forall l$, the value of $\zeta_{k}$ need to be set to
$\zeta_{k}=\frac{\nu_{\max,k}-r_{k}}{c},$ (81)
where $\nu_{\max,k}=\max\{\nu_{1}(\theta_{k},r_{k}),\dots,\nu_{N_{\mathrm{T}}}(\theta_{k},r_{k})\}$. Substituting (81) into (79) yields
$t_{k,l}=\frac{\nu_{\max,k}-\nu_{l}(\theta_{k},r_{k})}{c}.$ (82)
According to the geometric relationship illustrated in Fig. 4, the largest possible value of $t_{k,l}$ given in the above solution is obtained when $\theta_{k}=0$ or $\pi$, which is given by (c.f. (50))
$t_{k,l}\leq$ $\frac{\nu_{1}(\pi,r_{k})-\nu_{N_{\mathrm{T}}}(\pi,r_{k})}{c}=\frac{N(N_{\mathrm{T}}-1)d}{N_{\mathrm{T}}c}.$ (83)
Therefore, if $t_{\max}\geq\frac{N(N_{\mathrm{T}}-1)d}{N_{\mathrm{T}}c}$, the optimal solution in (82) can always be achieved. The proof is thus completed.